a test of a driver's perception reaction time is being conducted on a special testing track with wet pavement and a driving speed of 50 kph. when the driver is sober, a stop can be made just in time to avoid hitting an object that is visible 40 m. ahead. after a few drinks of san miguel beer, under exactly the same condition, the driver fails to stop in time and strikes the object at a speed of 30 kph. determine the driver's perception-reaction time after he was drinking. assume coefficient of friction is 0.60.


1
Expert's answer
2021-01-19T05:19:41-0500

"d= v"1 "\times tr"

"d=" distance travelled for perception

"v"1 = initial velocity

"tr"= perception time


"tr" ="\frac{d}{v1}" = "\frac{40}{50}= 0.8 seconds"

coefficient of friction "\mu" = "\frac{force applied }{normal weight }"= "\frac{ma}{mg}"


0.6="\frac{mg}{ma}= \frac{ma}{m\times 9.81}"

"a=9.81\times0.6=5.886 m/s"

"a" (acceleration/deceleration) = "\frac{final velocity- initial velocity}{time^2}"

= "\frac{mg}{t}" ="v"v- "v"u/t"2"

5.886= "\frac{0-30}{t^2}=equivalent = \frac{30}{t^2}"

t= "\sqrt[]{\frac{30}{5.886}} = 2.258 seconds"


total perception time for the unconscious driver is therefore prolonged as below;

t= 0.8 + 2.258

= 3.058 seconds